OpenStudy (anonymous):

How do I find the axis of symmetry for 3x^2+5x-4

jimthompson5910 (jim_thompson5910):

the axis of symmetry for y = ax^2 + bx + c is x = -b/(2a)

OpenStudy (anonymous):

How do I find the vertex?

jimthompson5910 (jim_thompson5910):

what did you get for the axis of symmetry?